OpenStudy (anonymous):

System of Linear equations: I added y=2x-8 to -3y+10=2 and I got -2y=2x+16.. I know I did something wrong- HELP!

OpenStudy (anonymous):

do you perhaps mean \[y=2x-8\] \[-3y+10x=2\]?

OpenStudy (anonymous):

if so replace y in the second equation by 2x-8. if not then i am not sure what the question is

OpenStudy (llort):

ok, we already know y, it is 8/3 we just need to find where x intersects that point it is 8/3=2x-8 isolate x from here...

OpenStudy (anonymous):

@ frownupsidedown what is the actual problem?

OpenStudy (llort):

he needs the point where the two lines intersect

OpenStudy (anonymous):

Solve the system of linear system equations by any method of choice y=2x-8 & -3y+10=2

OpenStudy (anonymous):

is the second equation \[-3y+10=2\] with no x in it?

OpenStudy (llort):

It is a horizontal line

OpenStudy (anonymous):

if so Llort is right

OpenStudy (anonymous):

Correct. That's probably why I'm so confused :P

OpenStudy (anonymous):

\[-3y+10=2\] \[-3y=-8\] \[y=\frac{8}{3}\] a horizontal line. then replace y by \[\frac{8}{3}\] in other equation to find x

OpenStudy (anonymous):

what Llor said. i thought maybe you had made a typo but second equation just fixes y at 8/3 and go from there

OpenStudy (anonymous):

@Llort sorry t doubt you

OpenStudy (llort):

not a problem

OpenStudy (anonymous):

Ohhh, Okay! Thank you so much! I get it now

OpenStudy (llort):

glad I could help